OpenStudy (anonymous):

Are the following equations parallel, perpendicular, the same line, or none of these 2x + 3y = -24 y = 2/3x - 24 A. Perpendicular B. Parallel C.None of these D. Same Line

OpenStudy (yellowlegoguy99):

A I think

OpenStudy (anonymous):

The answer is A, I graphed it.

OpenStudy (yellowlegoguy99):

I answered first, medal please? c:

OpenStudy (anonymous):

You just looked on my screen, cheater.

OpenStudy (yellowlegoguy99):

Like I understand graphs

OpenStudy (yellowlegoguy99):

On that cheating website, anyway

OpenStudy (texaschic101):

2x + 3y = -24 3y = -2x -24 y = -2/3x - 8 y = 2/3x - 24 slopes are -2/3 and 2/3 These are NEITHER parallel or perpendicular.

OpenStudy (anonymous):

Are you sure their not ??? (texaschic101)

OpenStudy (anonymous):

@texaschic101 Graph it here: https://www.desmos.com/calculator

OpenStudy (texaschic101):

for them to be parallel, the slopes have to be the same -- their not for them to be perpendicular, the slopes have to be negative reciprocals -- their not. I am sure....answer is C...none of these

OpenStudy (anonymous):

*they're

OpenStudy (texaschic101):

lol...this is why I am not in the english section

OpenStudy (texaschic101):

a perpendicular line will have negative reciprocal slopes. That means one slope would be 2/3 and the other would be -3/2. Which, they are not.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

I can confirm that both equation 1 has 2/3 for it's slope and equation 2 has -2/3 as its slope.

OpenStudy (texaschic101):

correct..... that means they are NEITHER parallel or perpendicular

OpenStudy (anonymous):

2x + 3y = -24 is equivalent to y = -2/3x - 8

OpenStudy (texaschic101):

yes

OpenStudy (anonymous):

y = -2/3x - 8 y = 2/3x - 24

OpenStudy (anonymous):

Oh, my bad. You're right. They're not at 90 degree angles.

OpenStudy (anonymous):

I forgot that part xD

OpenStudy (texaschic101):

for it to be perpendicular, one slope would have to be 2/3 and the other would have to be -3/2....

OpenStudy (anonymous):

Sorry guys but im going with the one whos right?

OpenStudy (texaschic101):

:)